### v1

# TACL: Threshold-Adaptive Curriculum Learning Strategy for Enhancing Medical Text Understanding

Abstract
Medical texts, particularly electronic medical records (EMRs), are a cornerstone of modern healthcare, capturing critical information about patient care, diagnoses, and treatments. These texts hold immense potential for advancing clinical decision-making and healthcare analytics. However, their unstructured nature, domain-specific language, and variability across contexts make automated understanding an intricate challenge. Despite the advancements in natural language processing, existing methods often treat all data as equally challenging, ignoring the inherent differences in complexity across clinical records. This oversight limits the ability of models to effectively generalize and perform well on rare or complex cases. In this paper, we present TACL (Threshold-Adaptive Curriculum Learning), a novel framework designed to address these challenges by rethinking how models interact with medical texts during training. Inspired by the principle of progressive learning, TACL dynamically adjusts the training process based on the complexity of individual samples. By categorizing data into difficulty levels and prioritizing simpler cases early in training, the model builds a strong foundation before tackling more complex records. By applying TACL to multilingual medical data, including English and Chinese clinical records, we observe significant improvements across diverse clinical tasks, including automatic ICD coding, readmission prediction and TCM syndrome differentiation. TACL not only enhances the performance of automated systems but also demonstrates the potential to unify approaches across disparate medical domains, paving the way for more accurate, scalable, and globally applicable medical text understanding solutions.
I Introduction
Medical texts, such as electronic medical records (EMRs), are vital for modern healthcare, capturing patient care, diagnoses, and treatments [1]. However, their unstructured and specialized nature poses significant challenges for automated understanding. EMRs often contain domain-specific terminology, inconsistencies, and varying complexity—from straightforward cases to intricate ones involving rare conditions or multiple comorbidities.
Figure 1: a) Illustration of Multi-Dimensional Complexity of Medical Text and Various Clinical Tasks relying on deep medical text understanding. b) A comparison of the performance of PLM-ICD on MIMIC-III-50-ICD9 without curriculum learning and with heuristic-based curriculum learning.
The complexity of EMRs spans multiple dimensions, as depicted in Figure 1(a). Narrative structure ranges from explicit and coherent records to fragmented and ambiguous texts requiring deeper reasoning. Clinical variability introduces further challenges, as common cases are easier to address, whereas rare or complex cases demand advanced understanding. Additionally, patient demographics contribute to variability, such as age-related differences in medical documentation. Traditional NLP models often often adopt a ”one-size-fits-all” paradigm, treating all medical texts as equally challenging [2, 3, 4, 5, 6, 7], leading to suboptimal performance, particularly on rare and complex cases that hold significant clinical importance.
This raises a key question: can we design a training paradigm that allows models to gradually and systematically acquire knowledge, just as humans learn? Recently, curriculum learning [8], a training scheme that aims to let the model learn from easier to more complex samples or tasks, can benefit machine learning algorithms. The efficacy of replacing conventional training with curriculum learning has been demonstrated in many application fields of healthcare, such as medical image segmentation [9, 10] and automatic ICD coding [5]. Inspired by this, we innovatively function medical text understanding as a progressive task with curriculum learning based on text complexity. However, two critical challenges remain: 1) how to design an effective difficulty assessment process for medical texts, and 2) how to develop a curriculum learning strategy that progressively guides the model from easier to harder examples. Intuitively, we can define the difficulty of medical text using some simple metrics like sequence length or code frequency. However, these strategies fail to improve performance and may even degrade it by overlooking semantic nuances, as shown in Figure 1(b). Therefore, how to determine a reliable and interpretable way to quantify the complexity of clinical narratives remains an urgent challenge. Additionally, how to design adaptive training strategies based on different tasks instead of manually setting thresholds, in order to ensure both task-specific performance improvements and the ability to generalize across diverse medical datasets remains an issue.
To address the two issues, we introduce TACL (Threshold-Adaptive Curriculum Learning) for medical text understanding, which addresses the challenges of quantifying text complexity and improving generalization across diverse medical datasets. To implement the curriculum strategy, we design a novel difficulty assessment method based on a domain-specific pre-trained language model, which effectively quantifies the complexity of clinical narratives by considering semantic, syntactic, and contextual features. Furthermore, we develop an adaptive training strategy that dynamically updates the curriculum to enhance the model’s ability to generalize across heterogeneous medical datasets.
TACL is validated on multilingual, multi-domain datasets, including English (MIMIC-III [11], MIMIC-IV [12]) and Chinese clinical records, demonstrating state-of-the-art performance in tasks like readmission prediction, out-hospital mortality, ICD coding, and TCM syndrome differentiation [13]. TACL achieves robustness in handling rare and complex cases, advancing the state of medical text understanding.
The main contributions can be summarized as follows:
  1. 1.
    To the best of our knowledge, this is the first work to apply curriculum learning to the medical text understanding task. By introducing curriculum learning into this domain, our approach provides a novel perspective for handling the inherent complexity and variability of clinical narratives.
  2. 2.
    We design a difficulty assessment method based on a domain-specific pre-trained language model, which effectively quantifies the complexity of medical texts. This method enables a systematic division of clinical narratives into difficulty levels, providing a structured foundation for the curriculum learning process.
  3. 3.
    We propose an adaptive training strategy that dynamically adjusts the learning process based on the model’s performance. This strategy enhances the model’s ability to generalize across diverse medical datasets, reducing overfitting to specific domains and improving robustness in real-world applications.
II Related Work
The challenges in automated medical text understanding have driven significant advancements in natural language processing (NLP) tailored for healthcare. This section reviews relevant work in two primary areas: curriculum learning and medical text understanding.
II-A Curriculum Learning in NLP
Curriculum learning, inspired by the human learning process, structures the training process by exposing models to simpler examples before progressively introducing more complex ones. Bengio et al.[8] first proposed this paradigm, which has shown success in various NLP tasks, including text classification, machine translation, and question answering. For example, Platanios et al.[14] introduced a dynamic curriculum framework for neural machine translation, where difficulty levels were determined by the model’s confidence.
While these methods have demonstrated efficacy, most rely on static or predefined difficulty measures, limiting their adaptability to the inherent variability of real-world data. Recent works on adaptive curriculum learning [15, 16] dynamically assess sample difficulty during training, allowing models to adjust to changing data distributions. However, these approaches primarily target general NLP tasks and fail to address the unique challenges of domain-specific data, such as medical texts. To bridge this gap, TACL introduces a threshold-adaptive curriculum mechanism tailored to the complexity of medical data. By dynamically prioritizing samples based on their difficulty, TACL improves both task-specific performance and generalization across diverse datasets.
II-B Medical Text Understanding
Automated medical text understanding plays a critical role in improving clinical decision-making and healthcare analytics. Early approaches relied on rule-based systems and traditional machine learning models, such as conditional random fields and support vector machines, which required extensive domain expertise and manual feature engineering. The rise of deep learning brought significant advancements, with CNNs and RNNs excelling in tasks like clinical text classification and named entity recognition.
Transformer-based models, such as BERT [17] and its domain-specific adaptations like ClinicalBERT [18] and BioBERT [19], have further advanced medical NLP by leveraging large-scale pretraining on biomedical corpora to capture domain-specific semantics. While effective, these models often treat all training samples equally, limiting their ability to handle rare or complex cases. TACL addresses this limitation by introducing a curriculum-based training strategy, enabling models to progressively learn from simpler to more complex cases and improving generalization.
Datasets like MIMIC-III [11] and MIMIC-IV [12] have been instrumental in advancing medical NLP, supporting tasks such as ICD coding and clinical risk prediction. However, the variability and complexity of these datasets highlight the need for adaptive frameworks like TACL to effectively handle heterogeneous and challenging data distributions.
Figure 2: Overview of our proposed TACL framework.
III Method
This section introduces our proposed framework, illustrated in Figure 2, which consists of four key stages: (1) generating contextual representations using domain-specific pre-trained models, (2) defining data difficulty levels via clustering, (3) dynamically adjusting the curriculum with the Threshold-Adaptive Curriculum Learning (TACL) strategy, and (4) applying task-specific prediction heads for downstream tasks. These steps progressively guide the model through increasingly complex patterns within the data, ensuring robust learning across diverse medical NLP challenges.
III-A Contextual Representation
To capture the semantic and terminological complexity of medical texts, we utilize domain-specific pre-trained models to generate high-dimensional contextual embeddings. Let the input dataset be denoted as:
$\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ (1)
where $x_{i}$ represents the input text, $y_{i}$ is the corresponding task-specific label (e.g., ICD code, readmission risk, or TCM syndrome), and $N$ is the total number of samples.
  • English Tasks: For tasks such as ICD coding out-of-hospital mortality, and readmission prediction, we employ RoBERTa [20] pre-trained on medical corpora.
  • Chinese Tasks: For TCM syndrome differentiation, we utilize ZY-BERT [13] pre-trained on traditional Chinese medicine datasets.
Thus, for each input text $x_{i}$, we extract the contextualized embedding of the [CLS] token, which serves as a summary representation of the entire input sequence. This embedding is obtained as follows:
$h_{i}=\text{Encoder}(x_{i})_{[CLS]},\quad h_{i}\in\mathbb{R}^{d}$ (2)
Here, $h_{i}$ represents the output embedding corresponding to the [CLS] token, which is widely used as the aggregate representation of the input sequence in transformer-based models. These embeddings form the foundation for subsequent clustering and curriculum learning steps.
III-B Data Clustering and Difficulty Definition
To implement curriculum learning, we define difficulty levels by clustering the contextual representations $\{h_{1},h_{2},\dots,h_{N}\}$. This process ensures that the curriculum is aligned with the intrinsic structure of the data.
Clustering
Given the embeddings $\{h_{i}\}_{i=1}^{N}$, we apply k-means clustering to partition the samples into $k$ clusters, $\mathcal{C}_{1},\mathcal{C}_{2},\dots,\mathcal{C}_{k}$. The objective of k-means clustering is to minimize the within-cluster variance (WCSS):
$\operatorname*{arg\,min}_{\mu_{1},\dots,\mu_{k}}\sum_{i=1}^{k}\sum_{h\in\mathcal{C}_{i}}\lVert h-\mu_{i}\rVert^{2}$ (3)
where $\mu_{i}\in\mathbb{R}^{d}$ is the centroid of cluster $\mathcal{C}_{i}$, defined as:
$\mu_{i}=\frac{1}{|\mathcal{C}_{i}|}\sum_{h\in\mathcal{C}_{i}}h$ (4)
Difficulty Definition
The difficulty of each cluster is determined by two factors:
  • Cluster Density: A measure of how tightly packed the embeddings are within a cluster. For cluster $\mathcal{C}_{i}$, the density is defined as:
    $\text{Density}(\mathcal{C}_{i})=\frac{1}{|\mathcal{C}_{i}|}\sum_{h\in\mathcal{C}_{i}}\lVert h-\mu_{i}\rVert^{2}$ (5)
    Lower density values indicate higher cluster density, meaning the samples are more homogeneous and less ambiguous.
  • Distance from Centroid: The average distance of samples in a cluster from the centroid. For cluster $\mathcal{C}_{i}$, it is defined as:
    $\text{Distance}(\mathcal{C}_{i})=\frac{1}{|\mathcal{C}_{i}|}\sum_{h\in\mathcal{C}_{i}}\lVert h-\mu_{i}\rVert$ (6)
    Lower distance values indicate that the samples are closer to the centroid, representing a smaller and less diverse cluster.
Clusters are classified into three difficulty levels based on their density and distance values: Easy clusters have low density values (high density) and low distance values, representing homogeneous and less diverse patterns; Medium clusters have moderate density and distance values, reflecting patterns with some complexity and diversity but not overly challenging; and Hard clusters have high density values (low density) and/or high distance values, representing highly diverse and ambiguous patterns.
III-C Threshold-Adaptive Curriculum Learning (TACL)
Traditional curriculum learning approaches often rely on static or heuristic criteria for transitioning between difficulty levels. In contrast, our TACL strategy dynamically adjusts the curriculum based on the model’s performance growth.
Growth Rate Metrics
Let the macro-F1 score at epoch $t$ be denoted as $F_{t}$. We define two growth rate metrics to monitor the model’s progress:
  • Average Growth Rate: Over a sliding window of the most recent $N$ epochs, the average growth rate is computed as:
    $\bar{\gamma}=\frac{1}{N-1}\sum_{t=2}^{N}\bigl(F_{t}-F_{t-1}\bigr)$ (7)
  • Instantaneous Growth Rate: The most recent growth rate is denoted as $\gamma_{\Delta}$ and is computed as:
    $\gamma_{\Delta}=F_{N}-F_{N-1}$ (8)
Transition Condition
A threshold coefficient $\beta\in(0,1)$ is introduced to determine when the model has sufficiently learned from the current difficulty level. The model is considered saturated at the current stage if:
$\gamma_{\Delta}<\beta\cdot\bar{\gamma}$ (9)
When saturation is detected, the training progresses to the next difficulty stage.
Curriculum Update
Let $\mathcal{D}_{\mathrm{easy}},\mathcal{D}_{\mathrm{middle}},\mathcal{D}_{\mathrm{hard}}$ represent the subsets of samples at each difficulty level. At each stage, training data is incrementally expanded:
[…]
At the final stage (hard), training continues until no improvement in macro-F1 is observed for $P$ consecutive epochs, after which training is terminated early.
The complete TACL framework is summarized in Algorithm 1.
Algorithm 1 Threshold-Adaptive Curriculum Learning
1:Input: Training data divided into difficulty levels $\mathcal{D}_{\mathrm{easy}},\mathcal{D}_{\mathrm{middle}},\mathcal{D}_{\mathrm{hard}}$, total epochs $E$, sliding window size NN, threshold coefficient $\beta$
2:Initialize: Current dataset $\mathcal{D}_{\mathrm{train}}=\mathcal{D}_{\mathrm{easy}}$, macro-F1 deque F=[]F=[\,] of length NN
3:for $t=1$ to $E$ do
4:  Train model on $\mathcal{D}_{\mathrm{train}}$ for one epoch
5:  Compute macro-F1 score $F_{\mathrm{current}}$
6:  Append $F_{\mathrm{current}}$ to $F$; discard oldest value if $|F|>N$
7:  if $|F|=N$ then
8:    Compute $\bar{\gamma}$ using equation (7)
9:    Compute $\gamma_{\Delta}$ using equation(8)
10:    if $\gamma_{\Delta}<\beta\cdot\bar{\gamma}$ then $\triangleright$ Check saturation
11:     if $\mathcal{D}_{\mathrm{train}}=\mathcal{D}_{\mathrm{easy}}$ then
12:      $\mathcal{D}_{\mathrm{train}}\leftarrow\mathcal{D}_{\mathrm{easy}}\cup\mathcal{D}_{\mathrm{middle}}$
13:     else if $\mathcal{D}_{\mathrm{train}}=\mathcal{D}_{\mathrm{easy}}\cup\mathcal{D}_{\mathrm{middle}}$ then
14:      $\mathcal{D}_{\mathrm{train}}\leftarrow\mathcal{D}_{\mathrm{easy}}\cup\mathcal{D}_{\mathrm{middle}}\cup\mathcal{D}_{\mathrm{hard}}$
15:     else
16:      Break: Early stop training if at hard stage
17:     end if
18:    end if
19:  end if
20:end for
21:Output: Trained model
III-D Downstream Task Prediction
After training with TACL, the model is fine-tuned on multiple downstream tasks using task-specific prediction heads, optimized with cross-entropy loss for classification tasks and binary cross-entropy loss for multi-label tasks.
TABLE I: Dataset statistics for the tasks. The training dataset is divided into three difficulty levels: Easy, Medium, and Hard.
Task Language Datasets Train Val Test
Easy Medium Hard Total
Task 1: Readmission Prediction English Discharge Summary 12,256 8,472 5,517 26,245 3,027 3,063
Early Notes 22,858 13,196 11,739 47,793 5,774 5,441
Task 2: Out-of-hospital Mortality English MIMIC-IV 14,437 12,823 6,227 33,487 4,186 4,186
Task 3: ICD Coding English MIMIC-III-full 33,024 12,963 1,732 47,719 1,631 3,372
MIMIC-III-50 3,437 2,389 2,240 8,066 1,573 1,729
MIMIC-IV-ICD10 49,158 38,142 1,789 89,089 13,378 19,799
Task 4: TCM Syndrome Differentiation Chinese TCM-SD 33,386 5,114 4,680 43,180 5,486 5,486
IV Experiment
IV-A Datasets
We evaluate our framework across four tasks using publicly available benchmark datasets, with each dataset partitioned into three difficulty levels (Easy, Medium, and Hard) described in Section 3.2. Table I provides an overview of dataset statistics for all tasks.
Task 1: Patient Readmission: We predict whether a patient will be readmitted to the hospital within a specific time frame based on clinical notes. Using the ADMISSIONS and NOTEEVENTS tables from the MIMIC-III dataset [11], two subsets are constructed:
  • Discharge Summary: Each sample consists of detailed discharge notes from NOTEEVENTS table, making it a rich source of information for prediction tasks.
  • Early Notes: Besides of discharge summary, it also provides a snapshot of a patient’s condition 3 days earlier in their care, offering complementary insights compared to discharge summaries.
Task 2: Out-of-hospital Mortality: Using discharge summaries from the MIMIC-IV dataset [12], this binary classification task predicts whether a patient dies within 30 days after discharge, excluding in-hospital deaths.
Task 3: ICD Coding: We evaluate our model on three datasets for ICD coding tasks, with each dataset representing either ICD-9 or ICD-10 coding:
  • MIMIC-III Full (ICD-9): Contains 52,723 clinical documents annotated with 8,929 unique ICD-9 codes. Each document has a median of 14 codes, making it a highly multi-label classification problem.
  • MIMIC-III 50 (ICD-9): Focuses on the top 50 most frequent ICD-9 codes, including 11,368 clinical documents. Each document is annotated with a median of 5 codes, representing a simpler but well-established benchmark.
  • MIMIC-IV (ICD-10): Provides clinical notes annotated with more granular ICD-10 codes, making this task more challenging due to the increased code specificity.
Preprocessing steps for all datasets include converting text to lowercase, removing non-alphanumeric characters, and truncating to a maximum of 2,500 or 4,000 tokens, depending on the model’s input size constraints.
Task 4: TCM Syndrome Differentiation: We classify Traditional Chinese Medicine (TCM) syndromes based on descriptions of patient conditions using the TCM-SD dataset [13]. Unlike the English-based tasks above, this dataset is in Chinese, with its unique linguistic structure and domain-specific terminology adding significant complexity.
TABLE II: Hyperparameter used for the tasks.
Task Model Hyperparameters
Batch Size Learning Rate Optimizer Epochs Length $\beta$
Task 1
Task 2
Bi-LSTM 64 0.0001 Adam 30 1024 0.8
Bio-BERT 64 0.00005 BertAdam 20 512 0.6
ClinicalBERT 64 0.00005 BertAdam 20 512 0.7
Task 3 Bi-GRU 8 0.001 AdamW 20 4000 0.7
CNN 8 0.001 AdamW 20 4000 0.7
CAML 8 0.005 AdamW 60 4000 0.5
MultiResCNN 16 0.0005 AdamW 40 4000 0.9
LAAT 8 0.001 AdamW 30 4000 0.7
PLM-ICD 8 0.00005 AdamW 20 4000 0.7
Task 4 RoBERTa 128 0.00005 AdamW 30 512 0.7
TCM-BERT 128 0.00005 AdamW 30 512 0.7
ZY-BERT 128 0.00005 AdamW 30 512 0.8
IV-B Implementation
Experiments are conducted using NVIDIA L40 48GB GPU. To ensure reproducibility, random seeds are fixed across all runs. The thresholds for growth rate in TACL are set from 0.5 to 0.8 according to the performance on validation subset, but may vary depending on the model. All relevant codes will be made publicly available [^1] https://anonymous.4open.science/r/medical-codes-354C.
For evaluating the performance of our proposed model, we employ three commonly used metrics: F1-score (Micro and Macro), AUC (Micro and Macro), and precision at K (P@K). The metrics are computed on the test set using the model that achieves the best performance on the validation set.
IV-C Baseline
We compare the proposed method against a range of strong baselines across all tasks. For tasks involving clinical and biomedical text, such as readmission prediction, out-of-hospital mortality, and ICD coding, we evaluate models like Bi-LSTM [21], which utilizes pre-trained Word2Vec embeddings and bidirectional LSTM layers, and transformer-based models like BioBERT [22] and ClinicalBERT [18], both pre-trained on large biomedical corpora. For ICD coding, additional specialized models like CAML [23] with label attention mechanisms and MultiResCNN [24] for multi-scale feature extraction are employed, alongside PLM-ICD [25], which integrates a pre-trained BERT encoder with a label attention mechanism. For TCM syndrome differentiation, we evaluate Chinese-language models like RoBERTa-wwm [26], pre-trained with Whole Word Masking on large-scale Chinese text, and ZY-BERT [13], a TCM-specific model trained on a large unlabelled corpus, achieving state-of-the-art performance on TCM tasks.
Figure 3: Performance comparison across four tasks: Patient Readmission (Task I), Out-of-hospital Mortality (Task II), ICD Coding (Task III), and TCM Syndrome Differentiation (Task IV).
V Main Results and Discussion
V-A Main Results
The performance of TACL is summarized across all four tasks in Figure 3, showcasing its consistent improvements over traditional training methods across various models and datasets. TACL enhances performance metrics such as Macro-F1, AUROC, and RP80, demonstrating its robustness and effectiveness in handling diverse medical text understanding challenges.
Notably, TACL provides significant gains for simpler models like Bi-GRU, CNN, and Bi-LSTM, particularly for tasks like readmission prediction and ICD coding, where these models often struggle with data complexity and rare labels. For example, TACL improves Macro-F1 scores and recall on rare codes, highlighting its ability to compensate for the limitations of simpler architectures.
TABLE III: Performance on Patient Readmission Prediction.
Model Stragegy Discharge Summary Early Notes
AUROC Accuracy AUPRC RP80 AUROC Accuracy AUPRC RP80
Logistic Regression Original 62.41 59.37 63.23 4.73 59.56 56.87 60.47 3.25
+Length 62.48 59.54 63.17 2.99 59.62 56.80 60.34 3.77
[…]
XGboost Original 60.14 57.59 63.48 2.12 58.40 56.76 60.78 2.04
+Length 60.33 57.70 63.51 3.18 59.00 56.93 61.18 1.79
[…]
SVM Original 63.26 58.67 63.21 6.59 58.49 56.03 61.17 3.16
+Length 63.07 58.59 63.18 5.55 59.01 56.57 61.39 4.98
[…]
TextCNN Original 63.79 59.56 66.71 4.65 60.15 57.06 62.37 6.46
+Length 63.65 59.49 66.53 3.10 60.11 56.99 62.16 5.56
[…]
Bi-LSTM Original 65.62 61.87 67.67 9.02 62.33 58.54 64.71 4.86
+Length 66.49 61.87 68.07 12.26 63.39 59.31 65.91 7.99
[…]
BERT Original 66.37 60.76 68.89 8.46 62.67 58.99 64.55 3.52
+Length 66.56 60.29 68.68 7.66 62.07 58.91 64.37 4.08
[…]
Bio-BERT Original 71.68 64.84 73.49 20.09 63.86 60.26 65.97 7.16
+Length 72.01 65.18 73.67 20.58 63.85 60.19 65.79 6.33
[…]
ClinicalBERT Original 75.34 68.76 75.83 27.17 64.80 60.93 66.64 8.82
+Length 73.73 68.46 74.81 23.34 64.04 59.77 66.24 10.67
[…]
Advanced models such as PLM-ICD, MultiResCNN, ClinicalBERT, and ZY-BERT also benefit from TACL, albeit with smaller but consistent gains. These improvements are especially pronounced in scenarios involving rare labels and imbalanced datasets, such as ICD-10 coding and TCM syndrome differentiation. TACL complements domain-specific pretraining by dynamically managing data complexity during training, leading to better generalization.
Overall, TACL’s consistent performance gains across both simpler and state-of-the-art models validate its adaptability and scalability for diverse tasks, including multilingual and domain-specific challenges. The results underscore its potential to enhance clinical decision-making and medical text understanding.
Figure 4: Ablation study and visualization about clustering number for readmission prediction task.
V-B Analysis on Difficulty Definition
The effectiveness of curriculum learning depends heavily on how data difficulty is defined. Figure 1(b) and Table III show that traditional statistical features including sequence length, number of code, and frequency of code, degrade performance compared to the baseline. These results indicate that statistical features fail to effectively capture the inherent complexity of clinical texts and may disrupt the learning process by introducing suboptimal sample ordering.
[…]
V-C Analysis on Data Order
[…]
This trend indicates that starting with simpler samples helps models build a solid foundation, enabling them to better handle more complex cases later in training. In contrast, reversing the order disrupts this progression, leading to suboptimal performance. These findings highlight the importance of a well-structured curriculum that aligns with the natural progression of model learning.
V-D Analysis on Clustering Number
The choice of clustering number directly impacts model performance and the interpretability of data distribution. As shown in the left panels of Figure 4, performance metrics across all three models and tasks achieve their peak when the clustering number is 3. This indicates that 3 clusters effectively balance the separation of data complexity levels, enabling models to learn optimally from progressively challenging samples.
The right panels of Figure 4 visualize the data distribution under different clustering numbers. When the clustering number is 3, the clusters are well-separated and intuitively aligned with the inherent structure of the data. Increasing the number of clusters beyond three results in over-segmentation, leading to redundant or overly granular clusters, while fewer clusters fail to capture sufficient complexity. These observations highlight the importance of selecting an optimal clustering number to achieve both strong performance and meaningful data partitioning.
V-E Clinical Applicability of TACL
The proposed TACL, demonstrates strong potential for clinical applications by addressing key challenges in medical text understanding tasks. As illustrated in the case study of ICD coding (Figure 5), predictions made without TACL often cluster around decision boundaries, leading to incorrect classifications with low confidence. By dynamically adjusting the learning process based on sample difficulty, TACL significantly improves the confidence of these borderline predictions, resulting in higher accuracy. This improvement not only enhances the overall performance of the model but also increases the trustworthiness of its predictions, a critical factor for clinical practitioners relying on AI-driven decision support systems, especially in high-stakes scenarios such as mortality prediction or rare disease diagnosis.
One of TACL’s key strengths lies in its adaptability, as it automatically learns sample difficulty and adjusts the curriculum to fit diverse tasks, datasets, and model architectures. Experiments across four medical text understanding tasks, involving both English and Chinese datasets, demonstrate consistent performance gains in metrics such as AUROC, Macro-F1, and AUPRC, highlighting its robustness in addressing challenges like imbalanced datasets, rare labels, and complex clinical narratives. From a clinical perspective, TACL’s ability to enhance prediction confidence and accuracy is particularly valuable in high-stakes scenarios, such as mortality prediction or rare disease diagnosis, where reliable and interpretable outputs are essential. By reducing manual fine-tuning efforts and improving trust in predictions, TACL provides a scalable solution for diverse medical applications. Furthermore, its principles can be extended to tasks such as multi-modal data integration or real-time predictions in electronic health records, making it a promising framework for advancing AI adoption in healthcare and improving patient outcomes.
Figure 5: Case study of ICD-9 coding
VI Conclusion
This paper introduces TACL, a novel framework designed to address the nuanced challenges of medical text understanding by dynamically adapting the training process to data complexity. Using contextual representation-based difficulty definitions and a curriculum learning strategy that progressively transitions from simpler to more complex samples, TACL achieves significant improvements in key medical NLP tasks, including ICD coding, readmission prediction, and differentiation of TCM syndrome. The results demonstrate TACL’s ability to enhance model generalization, particularly in handling rare and complex cases, which are often the most clinically significant. Furthermore, the framework’s adaptability to multilingual and domain-specific datasets, such as English clinical records and Chinese TCM texts, underscores its scalability and versatility. Beyond improving performance, TACL highlights the importance of aligning training strategies with task-specific nuances and data distributions, offering a transformative approach that bridges domain-specific challenges and scalable machine learning solutions.
Ethical Statement
The datasets used in this study are publicly available. Specifically, MIMIC-III and MIMIC-IV were accessed through PhysioNet after completing the required ethical training program. As such, this work adheres to ethical guidelines, and we do not foresee any ethical concerns arising from its use.